Solve by the elimination method.


4x-y=25
x+y=15

Answer by jim_thompson5910(35256)

Hint: add the two equations together (by adding the corresponding sides) to get

  4x-y=25
+  x+y=15
------------
  5x+0y=40


So we get the new equation 5x = 40. Solve this for 'x'. Once you have the solution, use it and plug it into any equation with 'y' to solve for 'y'.